OpenStudy (catearia):

Kenya exchanges $150 for British pounds(£). Suppose the conversion rate is £1 = $1.60. How many pounds should she receive?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (catearia):

have no clue

OpenStudy (skullpatrol):

What does " £1 = $1.60 " mean?

OpenStudy (smartperson):

first you would devide the amount of the conversion 1.60 by 150 dollars to get your answer

OpenStudy (smartperson):

oops multiply that

OpenStudy (skullpatrol):

try to set up the units to "cancel out"

OpenStudy (skullpatrol):

$$\Huge \dfrac{ £1}{$1.60}$$

OpenStudy (skullpatrol):

$$\Huge $150 \times \dfrac{ £1}{$1.60}$$

OpenStudy (mathmale):

skullpatrol is on target (as usual). " What does " £1 = $1.60 " mean?" This equation shows the equivalence of two different word currencies. If you bought something for 1 pound in London, you'd pay something like $1.60 back home for the same thing. If you have $150 and want to exchange that for British Pound Sterling, Make a conversion factor from £1 = $1.60: \[\frac{ £1 }{ $1.60 }=1\] I chose this because we want to eliminate $ in favor of £.

OpenStudy (mathmale):

Now multiply $150 by that conversion factor.